# Data-Juicer Sandbox: A Comprehensive Suite for Multimodal Data-Model Co-development

### v1

@@ Abstract
The emergence of large-scale multi-modal generative models has drastically advanced artificial intelligence, introducing unprecedented levels of performance and functionality. However, optimizing these models remains challenging due to historically isolated paths of model-centric and data-centric developments, leading to suboptimal outcomes and inefficient resource utilization. In response, we present a novel sandbox suite tailored for integrated data-model co-development. This sandbox provides a comprehensive experimental platform, enabling rapid iteration and insight-driven refinement of both data and models. Our proposed “Probe-Analyze-Refine” workflow, validated through applications on state-of-the-art LLaVA-like and DiT-based models, yields significant performance boosts, such as topping the VBench leaderboard. We also uncover fruitful insights gleaned from exhaustive benchmarks, shedding light on the critical interplay between data quality, diversity, and model behavior. With the hope of fostering deeper understanding and future progress in multi-modal data and generative modeling, our codes, datasets, and models are maintained and accessible at https://github.com/modelscope/data-juicer/blob/main/docs/Sandbox.md.
@@ 1 Introduction
The advent of multi-modal generative models has revolutionized artificial intelligence, pushing the boundaries of functionality and creativity across various domains (OpenAI, 2024a; b; Wang et al., 2024). Recognizing the pivotal role of training data in shaping model performance, there are fast-growing efforts to curate datasets of larger scales and higher quality (Jakubik et al., 2024).
@@
However, the development trajectories of these models and datasets have historically diverged, guided more by intuition than by systematic co-development methodologies. Recent advances in enhancing multi-modal generative models tend to be either model-centric or data-centric, rarely bridging the two aspects cohesively. For example, model-centric methods focus on algorithmic enhancements and architectural innovations under fixed data priors, while data-centric strategies usually concentrate on processing and cleaning datasets independently of specific model training contexts (Qin et al., 2024). Both approaches usually suffer from a lack of systematic guidance and cooperative synergy, relying heavily on heuristic exploration and single-perspective expertise. This fragmented landscape presents a significant barrier to achieving optimal model performance, as the interplay between data characteristics and model capabilities remains largely underexploited.
@@
Moreover, the practical implementation of multi-modal generative models is further complicated by infrastructure constraints, escalating computational costs, and the accelerating pace of development cycles (Xu et al., 2024b). In the age of large-scale models with rapidly growing model parameters and dataset sizes, the processes of data processing and model training become increasingly resource-intensive, demanding substantial time and computations. Due to the absence of cost-effective platforms that simplify and speed up data-model co-development, researchers and developers often face the dilemma of prioritizing result-driven development at the expense of thorough, insight-led exploration. This deficiency hinders the iterative refinement for both domains, leading to sub-optimal outcomes as improvements in one domain are hard to inform, apply and enhance each other directly.
@@
To fill this gap, we introduce the Data-Juicer Sandbox, a comprehensive suite for facilitating the co-development of multi-modal data and generative models. Building upon an open-source data processing system tailored for multi-modal generative models, Data-Juicer (Chen et al., 2024a), our sandbox suite further integrates a wealth of off-the-shelf components optimized for usability and compatibility with existing model-centric infrastructures. Collectively, it offers flexibly customizable orchestration from different levels including end-to-end workflows, specific development behaviors, and underlying data-model development capabilities. Within the sandbox laboratory, users are empowered to rapidly explore different data pipelines and model configurations on cost-controlled datasets and models. This accelerates the discovery of insightful patterns and informed decision-making, ultimately paving the way for scalable, resource-efficient data-model development.
@@
To exemplify the efficacy of our sandbox, we propose a “Probe-Analyze-Refine” workflow, meticulously crafted to explore the synergies between data processing operators, target model metrics, and the scalability of these enhancements. We apply this workflow to two cutting-edge models: Mini-Gemini (Li et al., 2023), an LLaVA-inspired model for image-to-text generation, and EasyAnimate (Xu et al., 2024a), a Diffusion Transformer based model for text-to-video generation. Thanks to the sandbox’s capabilities, we attain significant advancements in data quality and model performance, such as achieving top spot on the VBench (Huang et al., 2024) leaderboard, outperforming strong competitors such as VideoCrafter (Chen et al., 2024b) and AnimateDiff (Guo et al., 2024). These achievements are underpinned by a series of insights linking 31 data processing operators and 35 model metrics, including analysis of the fine-grained impact of data processing for model training, the delicate balance between data diversity and model performance, and the strategic optimization of data scaling for enhanced model-data co-development.
@@
  • Innovative Sandbox Suite: To the best of our knowledge, this is the first open-source sandbox suite tailored for co-development between multi-modal data and generative models, rendering experimental exploration in this field more insightful, systematic, convenient and reusable.
  • Effect-Proven Workflow: We present a new progressive workflow for data-model co-development and substantiate its impact through extensive empirical evidence such as achieving new top-tier performance in image understanding and video generation tasks.
  • Practical Guidance: We conduct extensive experiments on benchmarking the effects of dozens of data processing operators and model metrics, providing fruitful and valuable insights toward further advancements in multi-modal generative models.
@@ 2 Related Works
Model-Centric Progress in Multimodal Generative Models.
Multimodal generative models have captivated researchers with their formidable capabilities (OpenAI, 2024a; b), leading to a surge in model-centric development efforts. These focuses mainly lie in refining training algorithms (Caffagni et al., 2024; Li et al., 2024; Zhang et al., 2024), advancing model architectures and components (He et al., 2024a; Yin et al., 2024; Jiao et al., 2024), and harnessing the models’ generative potential for various applications (Wang et al., 2024; Liu et al., 2024a; Zhou et al., 2024). There is a growing consensus that transformer-based scaling is predominant (Xu et al., 2023). However, the high computational requirements imposed by scaling laws (Xu et al., 2024b) and the optimization challenges inherent to generative models (Manduchi et al., 2024) often confine insights to specific datasets or vague data characteristics. This situation leaves a significant gap in comprehending the extent to which models’ performance and behavior hinge upon implicit assumptions and inductive biases embedded within the underlying data distributions.
Trends in Data-Centric Development for Multimodal Generative Models.
An emerging trend shifts the focal point from models to data (Jakubik et al., 2024; Bai et al., 2024), underscored by the notion that large models function akin to data compressors (Delétang et al., 2024). Echoing the principle of “garbage in, garbage out”, meticulous data processing is recognized as pivotal. Efforts now isolate data manipulation as a primary experimental variable in multimodal generative modeling (He et al., 2024b). Nonetheless, multimodal data processing involves highly heterogeneous processing workflows, vast quantities, diverse types, and the high cost of training downstream models. This complexity results in predominantly heuristic approaches, such as data filtering and synthesis guided by human intuition (Long et al., 2024). There remains a huge exploration space for more systematic data development methodologies and their harmonious integration with model development.
Open-Source Infrastructure for Multimodal Generative Model Development.
The ecosystem has flourished with robust infrastructure for multimodal model development, including frameworks for training and evaluation like Transformers (Wolf et al., 2020), Diffusers (von Platen et al., 2022), NeMo (Harper et al., ), MMagic (MMagic Contributors, 2023), ESPNet (Peng et al., 2023), MMBench (Liu et al., 2024b) and etc. Regarding multimodal data infrastructure, contributions are mainly in the form of datasets and dataset-specific preprocessing scripts, such as DatasetsHub from HuggingFace (Lhoest et al., 2021). The practical wisdom and foundational capabilities in data processing often remain unstandardized and underleveraged. Existing data processing frameworks, mainly tailored for single-modal data (Computer, 2023; Bradski, 2000; Hwang et al., 2023), highlight the nascent stage of systematic multimodal data development platforms (Chen et al., 2024a; Du et al., 2023). Recognizing the reciprocal relationship where large, high-quality datasets facilitate superior models, which in turn can aid in generating more refined datasets, our work introduces a dedicated intermediary layer. By seamlessly integrating state-of-the-art model-centric multimodal infrastructure with the Data-Juicer system, we enable streamlined, insight-driven co-development between models and data.
3 The Proposed Data-Juicer Sandbox Laboratory
3.1 Motivation and Design
Figure 1: Overview of the Data-Juicer Sandbox Laboratory. Here, Probe Jobs conduct preliminary analyzes on both data and models. Based on the insights gleaned from probe results, Refine-Recipe Jobs optimize data processing recipes and Execution Jobs subsequently undertake the core tasks of data processing, model training, and inference. Lastly, Evaluation Jobs serve to critically appraise the resultant data quality and model effectiveness. These four stages proceed in a methodical sequence, with each step offering versatile customization through configuration options at diverse granularity across the data-model co-development lifecycle: end-to-end workflows, generic behaviors, and foundational capabilities.
Why do we need data-model co-development?
@@ 3.1 Motivation and Design
In the era of large models, the development of both data and models necessitates a collaborative approach involving numerous algorithm researchers and system engineers. Training data for large models is often highly heterogeneous in terms of quality, context, type, and timeliness. The processing and mixing (Ge et al., 2024) of these datasets, known as data recipes, are complex and varied. The scale of data amplifies the stakes for refinement attempts on both data and models, imposing substantial computational and time burdens. Traditional data-centric or model-centric strategies fall short by optimizing in isolation, leading to diminished overall efficiency and resource misallocation. When either component requires adjustment, the dual optimization challenge inflates costs, as one part may have already reached near-optimal status.
Why do we need a sandbox laboratory?

### v2

# Data-Juicer Sandbox: A Feedback-Driven Suite for Multimodal Data-Model Co-development

@@ Abstract
The emergence of multimodal large models has advanced artificial intelligence, introducing unprecedented levels of performance and functionality. However, optimizing these models remains challenging due to historically isolated paths of model-centric and data-centric developments, leading to suboptimal outcomes and inefficient resource utilization. In response, we present a new sandbox suite tailored for integrated data-model co-development. This sandbox provides a feedback-driven experimental platform, enabling cost-effective iteration and guided refinement of both data and models. Our proposed “Probe-Analyze-Refine” workflow, validated through practical use cases on multimodal tasks such as image-text pre-training with CLIP, image-to-text generation with LLaVA-like models, and text-to-video generation with DiT-based models, yields transferable and notable performance boosts, such as topping the VBench leaderboard. Extensive experiments also uncover fruitful insights into the interplay between data quality, diversity, model behavior, and computational costs. All codes, datasets, and models are open-sourced to foster future research and applications that would otherwise be infeasible due to the lack of a dedicated co-development infrastructure.
Machine Learning, ICML
@@ 1 Introduction
The advent of multimodal large models has revolutionized artificial intelligence, pushing the boundaries of functionality and creativity across various domains (OpenAI, 2024a; Wang et al., 2024a). Recognizing the key role of data in shaping model performance, there are fast-growing efforts to curate datasets of larger scales and higher quality (Jakubik et al., 2024; Gadre et al., 2023; Xu et al., 2024a).
@@
However, the development trajectories of these models and datasets have historically diverged, guided more by intuition than by systematic co-development methodologies. Recent advances in enhancing multimodal large models tend to be either model-centric or data-centric, rarely bridging these two aspects closely. For example, model-centric methods focus on algorithmic enhancements and architectural innovations under fixed data priors, while data-centric strategies usually focus on processing datasets independently of specific model training contexts (Qin et al., 2024). Both approaches usually suffer from a lack of systematic principles and cooperative synergy, relying heavily on heuristic exploration and single-perspective expertise. This fragmented landscape presents a notable barrier to achieving optimal performance, as the interplay between data characteristics and model capabilities remains largely under-exploited.
@@
Moreover, putting multimodal large models into practice is further complicated by infrastructure limitations, increasing computational costs, and the faster pace of the project development (Xu et al., 2024c). In the age of large models with rapidly growing model parameters and dataset sizes, the processes of data processing and model training become increasingly resource-intensive, demanding substantial time and computations. Due to the absence of cost-effective platforms that simplify and speed up data-model co-development, researchers and developers often face the dilemma of prioritizing result-driven development at the expense of thorough, insight-led exploration. This deficiency hinders the iterative refinement for both domains, leading to sub-optimal outcomes as improvements in one domain are hard to inform, apply and enhance each other directly.
@@
To fill this gap, we introduce the Data-Juicer Sandbox, a feedback-driven suite for facilitating the co-development of multimodal data and models. Building upon an open-source data processing system tailored for multimodal large models, Data-Juicer (Chen et al., 2024a, b), our sandbox suite integrates a wealth of off-the-shelf components optimized for usability and compatibility with open-source model-centric infrastructures. Collectively, it offers customizable orchestration from different levels including end-to-end workflows, specific development behaviors, and underlying data-model development capabilities. Within the sandbox laboratory, users are empowered to optimize data and models based on their fruitful feedback in cost-controlled environment. This accelerates insight discovery and informed decision-making, paving the way for transferable, resource-efficient data-model optimization.
@@
To exemplify the efficacy of the sandbox, we propose a “Probe-Analyze-Refine” workflow, crafted to explore the interplay between data processing operators (OPs), model performance feedback, and the transferability of these enhancements. We apply this workflow to four cutting-edge models: Mini-Gemini (Li et al., 2024b), an LLaVA-inspired model for image-to-text generation, EasyAnimate (Xu et al., 2024b) and T2V-Turbo (Li et al., 2024a), two Diffusion Transformer based models for text-to-video generation, and a CLIP model (Gadre et al., 2023) for image-text pre-training. Our extensive experiments show notable evidence of the Sandbox’s usability, including topping the VBench (Huang et al., 2024) leaderboard with superior performance than competitors such as Gen-3 (RunwayML, 2024) and VEnhancer (He et al., 2024a). The improvement is underpinned by a series of insights linking more than 40 data processing OPs and 70 benchmark metrics, with fine-grained analysis on the balance between data quality, diversity, compute cost, and model performance.
@@
  • To the best of our knowledge, this is the first open-source  [^1] https://github.com/modelscope/data-juicer/blob/main/docs/Sandbox.md sandbox suite tailored for co-development between multimodal data and models, rendering experimental exploration in this field more insightful and systematic.
  • We present a new effect-proven workflow for data-model co-development and substantiate its impact through empirical evidence among image understanding, video generation, and image-text pretraining tasks.
  • We conduct extensive experiments on benchmarking the effects of dozens of data processing operators and model metrics, providing practical guidance toward further advancements in multimodal large models.
@@ 2 Related Works
Model-Centric Progress in Multimodal Large Models. Multimodal large models have gained prominence for their remarkable capabilities (OpenAI, 2024a, b). Advances in training algorithms (Caffagni et al., 2024; Li et al., 2024a) and model architectures (He et al., 2024a; Yin et al., 2024) have fueled this interest. Transformer scaling remains a prevalent approach (Xu et al., 2023), though high computational demands and optimization challenges often restrict insights to specific datasets and create a gap in understanding how implicit data biases affect model performance.
Trends in Data-Centric Development. Recently, a shift towards data-centric development has emerged (Jakubik et al., 2024), emphasizing data handling as key to the efficacy of large models such as CLIP (Gadre et al., 2023). Despite the increasing recognition of data processing, the heterogeneous nature of multimodal data leads to predominantly heuristic approaches (Long et al., 2024), underscoring the pressing need for more systematic methodologies for data-model co-development.
Open-Source Infrastructure. The ecosystem for multimodal model development has expanded with fruitful open-sourced frameworks (Wolf et al., 2020; Liu et al., 2023). However, contributions to multimodal data infrastructure often consist of raw datasets and preprocessing scripts, lacking standardized practices. Most existing data processing frameworks are tailored for single-modal data (TogetherAI, 2023), highlighting the early stage of multimodal data development. To address these gaps, our work presents a new intermediary layer that connects advanced model infrastructures with the Data-Juicer system, facilitating better co-development between models and data.
Detailed comparison on more works are in Appendix A.
Figure 1: A probe-analyze-refine workflow within Data-Juicer sandbox for systematic data-model co-development. (1) The process begins with cost-effective, contrastive explorations on small data pools to identify the most effective data processing operators (OPs). (2) These OPs are then combined and scaled within data recipes through hierarchical data pyramids to assess their cumulative impact on model performance. (3) Lastly, data utilization is optimized via duplication and diversity analysis, guiding the final formulation of refined data recipes and datasets for enhanced large-scale training. This structured workflow enables informed decision-making and efficient optimization for developing superior models.
3 The Proposed Sandbox Laboratory
3.1 Motivation and Overview
@@
Why do we need data-model co-development? In the era of large models, the development of both data and models necessitates collaboration involving numerous algorithm researchers and system engineers. Training data for large models is often highly heterogeneous in terms of quality, context, type, and timeliness. The processing and mixing of these datasets, known as data recipes, are complex and varied (Ge et al., 2024). The scale of data amplifies the stakes for refinement attempts on both data and models, imposing substantial computational and time burdens. Traditional data-centric or model-centric strategies fall short by optimizing in isolation, leading to diminished overall efficiency and resource misallocation (Qin et al., 2024). When either component requires adjustment, the dual optimization challenge inflates costs, as one part may have already reached near-optimal status.